# Computational and Nonlinear Dynamics

### v1

@@
[…]
Computational and Nonlinear Dynamics
[…]
Auburn University,
Auburn, AL 36849,
email : mzm0390@auburn.edu
[…]
Dan B. Marghitu Department of Mechanical Engineering,
Auburn University,
Auburn, AL 36849,
email : marghdb@auburn.edu
Asymptotic Behavior of an Unforced Duhem-Type Hysteretic Oscillator
(5 janvier 2026)
Résumé
@@ 1 Introduction
The Duhem models constitute a class of differential models of hysteresis that are suitable for the description of a variety of physical phenomena, including dry friction, elastoplastic materials, and magnetization. The models are usually attributed to Pierre Duhem [Duhem (1896-1902), as cited in ikhouane_survey_2018]. However, the interest in the models has surged only in the second half of the twentieth century [everett_general_1954, chua_mathematical_1971, chua_generalized_1972, pokrovskii_theory_1973, visintin_continuity_1983, coleman_constitutive_1986, coleman_class_1987, krasnoselskii_systems_1989, visintin_differential_1994, oh_modeling_2003, oh_analysis_2003, oh_semilinear_2005, padthe_counterclockwise_2005, jayawardhana_sufficient_2009, jayawardhana_dissipativity_2011, jayawardhana_stability_2012, ouyang_stability_2012, naser_consistency_2013, naser_characterization_2013, vasquez-beltran_modeling_2023]. Specializations of the models include the Dahl friction model [dahl_solid_1968], the LuGre friction model [canudas_de_wit_new_1995], and the Bouc-Wen model [bouc_forced_1967, bouc_modemathematique_1971, wen_method_1976]. Refs. [visintin_differential_1994] and [ikhouane_survey_2018, ikhouane_erratum_2018] provide further general information about the Duhem models. Refs. [krasnoselskii_systems_1989, mayergoyz_mathematical_1991, visintin_differential_1994, mielke_rate-independent_2015] provide further information about differential models of hysteresis in general.
The following form of a Duhem model is essentially equivalent [^1] Apart from syntactic discrepancies, Ref. [oh_analysis_2003] employs a different set of regularity conditions for $f$, $g$, and $h$. to the form employed in Ref. [oh_analysis_2003] : [^2] See Appendix A for notation and conventions.
@@
The article addresses the problem of the asymptotic behavior of solutions of the system given by Eq. (8). More specifically, it is shown that every trajectory of the system converges to an equilibrium point. This result is a generalization of a result about the asymptotic behavior of the Bouc-Wen model of class I that was previously presented in Refs. [ikhouane_dynamic_2007, ikhouane_systems_2007]. It should be noted that other closely related problems were considered in [anderssen_global_1998, jayawardhana_stability_2012, ouyang_stability_2012, ouyang_absolute_2014].
@@ 4 The Bouc-Wen Oscillator
which is equivalent to Eq. (8). It can be verified that the functions $f_{1}$, $f_{2}$, $g_{1}$, $g_{2}$, $h_{1}$, $h_{2}$, $c$ satisfy the conditions necessary for the application of Theorem 3.14 provided that $A\in\mathbb{R}_{>0}$, $\gamma\in(-\beta,\beta]$, and $b\in\mathbb{R}_{\geq 0}$ (the Bouc-Wen model of class I). In this case, by Theorem 3.14, there exist $x_{\infty},z_{\infty}\in\mathbb{R}$ such that $\lim_{t\rightarrow+\infty}x(t)=x_{\infty}$, $\lim_{t\rightarrow+\infty}z(t)=z_{\infty}$, $\lim_{t\rightarrow+\infty}v(t)=0$, and $\alpha x_{\infty}+(1-\alpha)Az_{\infty}=0$.
@@ 5 Conclusions and Future Work
Future work may include generalization of the results presented in this article to other (more abstract) Duhem-type models, an investigation of the stability of individual equilibrium points, and estimation of the rate of convergence of the solutions.
@@ Acknowledgment
Annexe A Notation, Conventions, Foundations

### v2

@@
Abstract
@@ 1 Introduction
The Duhem models constitute a class of differential models of hysteresis that are suitable for the description of a variety of physical phenomena, including dry friction, elastoplastic materials, and magnetization. The models are usually attributed to Pierre Duhem [Duhem (1896-1902), as cited in ikhouane_survey_2018]. However, the interest in these models has surged only in the second half of the twentieth century [everett_general_1954, chua_mathematical_1971, chua_generalized_1972, pokrovskii_theory_1973, visintin_continuity_1983, coleman_constitutive_1986, coleman_class_1987, krasnoselskii_systems_1989, visintin_differential_1994, oh_modeling_2003, oh_analysis_2003, oh_semilinear_2005, padthe_counterclockwise_2005, jayawardhana_sufficient_2009, jayawardhana_dissipativity_2011, jayawardhana_stability_2012, ouyang_stability_2012, naser_consistency_2013, naser_characterization_2013, vasquez-beltran_modeling_2023]. Specializations of the models include the Dahl friction model [dahl_solid_1968], the LuGre friction model [canudas_de_wit_new_1995], and the Bouc-Wen model [bouc_forced_1967, bouc_modemathematique_1971, wen_method_1976]. Refs. [visintin_differential_1994] and [ikhouane_survey_2018, ikhouane_erratum_2018] provide further general information about the Duhem models. Refs. [krasnoselskii_systems_1989, mayergoyz_mathematical_1991, visintin_differential_1994, brokate_hysteresis_1996, mielke_rate-independent_2015] provide further information about differential models of hysteresis.
The following form of a Duhem model is essentially equivalent [^1] Apart from syntactic discrepancies, a different set of regularity conditions for $f$, $g$, and $h$ is employed in Ref. [oh_analysis_2003]. to the form employed in Ref. [oh_analysis_2003]: [^2] See Appendix A for notation and conventions.
@@
The article addresses the problem of the asymptotic behavior of solutions of the system given by Eq. (8). More specifically, it is shown that every trajectory of the system converges to an equilibrium point. This result is a generalization of a result about the asymptotic behavior of the Bouc-Wen model of class I that was previously presented in Refs. [ikhouane_dynamic_2007, ikhouane_systems_2007]. It should be noted that other related problems were considered in [yakubovich_method_1965, barabanov_absolute_1979, anderssen_global_1998, lacy_hysteretic_2000, logemann_systems_2003, logemann_class_2008, jayawardhana_stability_2012, ouyang_stability_2012, ouyang_absolute_2014].
@@ 4 The Bouc-Wen Oscillator
which is equivalent to Eq. (8). It can be verified that the functions $f_{1}$, $f_{2}$, $g_{1}$, $g_{2}$, $h_{1}$, $h_{2}$, $c$ satisfy the conditions necessary for the application of Theorem 3.14 provided that $A\in\mathbb{R}_{>0}$, $\gamma\in(-\beta,\beta]$, and $b\in\mathbb{R}_{\geq 0}$ (the Bouc-Wen model of class I). In this case, by Theorem 3.14, there exist $x_{\infty},z_{\infty}\in\mathbb{R}$ such that $\lim_{t\rightarrow+\infty}x(t)=x_{\infty}$, $\lim_{t\rightarrow+\infty}z(t)=z_{\infty}$, $\lim_{t\rightarrow+\infty}v(t)=0$, and $\alpha x_{\infty}+(1-\alpha)Az_{\infty}=0$. This is equivalent to $\lim_{t\rightarrow+\infty}x(t)=x_{\infty}$, $\lim_{t\rightarrow+\infty}z(t)=z_{\infty}$, $\lim_{t\rightarrow+\infty}v(t)=0$, and $\alpha x_{\infty}+(1-\alpha)Dz_{\infty}=0$ in the original coordinates (under the $z_{\infty}\mapsto AD^{-1}z_{\infty}$ transformation).
@@ 5 Conclusions and Future Work
Future work may include generalization of the results presented in this article to other (more abstract) Duhem-type models, an investigation of the stability of individual equilibrium points, estimation of the rate of convergence of the solutions, and investigation of the behavior of the system under the influence of external disturbances.
@@ Acknowledgment
Appendix A Notation, Conventions, Foundations